jaynater:

If a trapezoid has bases of 15 and 12, and has an area of 67.5, what is the height?

Aureo:

All you have to do here is divide 67.5 by 15 first, then divide that number by 12.

Aureo:

So what do you think the answer is?

Aureo:

If you need help just ask me

jaynater:

By the way, the formula for a trapezoid is "Area=1/2 height(base 1+base 2)"

Ultrilliam:

^That is how to get the area, not how to get the height from the area =P

jaynater:

Yes, but you'll need to find the height to get the area

jaynater:

which is a given area

Ultrilliam:

Although, looking at it now, he was indeed wrong. You would calculate it like this \(H=2\times[a\div(B1+B2)]\) H=2*(Area/(Base1+Base2)) So in this instance it would have been H=2*(67.5/(15+12))

jaynater:

I had it backwards, but still got the answer right...

jaynater:

Thanks btw

Ultrilliam:

You're welcome!